application of linear equations
Andrew invested a total of Php18,000, part at 5% and part at 9%. If his income for one year from the 9% investment was Php100 less than his income from the 5% investment, how much was invested at each rate?

Answer by mananth(16946)

Andrew invested a total of 18,000,
part at 5% . Let this investment be x
interest = 0.05x
..
part at 9%. this will be 18000-x
interest will be (18000-x )*0.09
..
interest from the 9% investment was Php100 less than his income from the 5%
..
income from 5% - income from 9% = 100
..
0.05x - 0.09(18000-x)= 100
0.05x-1620+0.09x=100
0.14x=1620+100
0.14x=1720
x= 1720/0.14
= Php 12,286
..
investment at 9% will be 18000-12286
= Php 5,714
CHECK(0.05*12286) - 0.09(18000-12286)=100